princess765:

a scanner scanned 72 photos in eight minutes. if it scans photos at a constant rate, it can scan _____photes in 23 minutes

mxddi3:

Welcome to QC!! Since we have 72 photos in 8 minutes, but we need to find out how many are going to be scanned in 23 minutes, we are going to need to find the unit rate first. \[\frac{ 72 }{ 8 }= \frac{ ? }{ 1 }\] in order to get from 8 to 1, we have to divide by 8. since we divided by 8 on the bottom, we are going to do the same for the top. \[72\div8= 9 \] this means our new UNIT RATE is 9 photos scanned every ONE minute. now, since we need to know how many are scanned in 23 minutes, we can do: \[\frac{ 9 }{ 1 }=\frac{ ? }{ 23 }\] to get from one minute to 23 minutes, we will MULTIPLY the bottom by 23. since we multiplied the bottom by 23, we are going to multiply the TOP by 23. so, \[9\times23=?\]